OpenStudy (anonymous):

Find the slope of line passing through the points (5, 9) and (5, -2) A. -11 B. 0 C. 11 D. undefined

OpenStudy (anonymous):

@luckythebest check my work okay :)

OpenStudy (anonymous):

give the answer; i'll see

OpenStudy (anonymous):

C

OpenStudy (anonymous):

9-(-2)=11 5-5=o so its C

OpenStudy (anonymous):

Yea ^

OpenStudy (anonymous):

@ilovedaivd what is 11/0, as far as i know; anything divided by 0 is undefined.

OpenStudy (anonymous):

@DaWizjr anything divided by 0 = undefined. D

OpenStudy (anonymous):

C

OpenStudy (anonymous):

:o oooo i never new that X)

OpenStudy (e.cociuba):

@luckythebest is correct.

OpenStudy (anonymous):

It is C

OpenStudy (anonymous):

NO its C

OpenStudy (anonymous):

Im telling you its C

OpenStudy (anonymous):

Ok dont listen to me and fail

OpenStudy (e.cociuba):

No cuz 11/0= undefined @DaWizjr

OpenStudy (anonymous):

@ilovedaivd and if you carefully saw that the x-coordinate is the same. Also remember this; slope of a line is also the Tan x of the line with angle x in positive direction of x-axis. Since x = 5 is a perfectly straight line, the angle made by this line with x-axis is 90. And Tan 90 is not defined. I'll explain it with a diagram. wait.

OpenStudy (e.cociuba):

Back up ur answer if ur so sure hun, but don't go messing up what others know is correct @DaWizjr

OpenStudy (anonymous):